Youve observed the following returns on SkyNet Data Corporations stock over the past five years: 20 percent, – 12 percent,
Answer #1

Average return=Total return/Total time period

=(20-12+17+20+10)/5=11%

Return (Return-Average Return)^2
20 (20-11)^2=81
-12 (-12-11)^2=529
17 (17-11)^2=36
20 (20-11)^2=81
10 (10-11)^2=1
Total=728%

Standard deviation=[Total (Return-Average Return)^2/(Time period-1)]^(1/2)

=[728/(5-1)]^(1/2)

=[728/4]^(1/2)

=13.49%(Approx)

Variance=Standard deviation^2

=0.0182